Answer question 4
0 Convenience sampling QUESTION 4 Lets say there are 10,000 lawyers in the USA and 500 of them are Oreo cookie lovers. These 500 lawyers consume a total of 500 Oreo cookies in a given time period out of 2,000 cookies sold. What is the BDI for Oreo cookies consumed by lawyers 100 Θ50 1000 500 QUESTION 5 If your measured responses to survey questions over two different points of time were drastically different, we would say that this question isnt meeting the criteria of
Answer #1

BDI = (% Market Brand Sales / % Market Population) x 100

% Market Brand sales=500/2000=0.25

%Market population=500/10000=0.05

Hence BDI=0.25/0.05*100=500

Hence last option is correct